# Experimental validation of a linear momentum and bluff-body model for high-blockage cross-flow turbine arrays

### v1

@@ 4.4 Bluff-Body Scaling of Performance Data
Figure 17: Time-averaged array performance from figure 6 scaled by 17 the corresponding analytical bypass velocities (figure 12) and 17 the unconfined freestream velocity calculated using Eq. 19.
Given that open-channel LMADT is descriptive of the measured bypass flow across a range of array operating conditions, we turn to the question of whether LMADT can also be descriptive of the measured array performance. In prior work, LMADT has been most commonly used in analytical blockage corrections to estimate unconfined performance from measurements or simulations of confined operation (e.g., Bahaj et al., 2007; Chen & Liou, 2011; Kinsey & Dumas, 2017; Ross & Polagye, 2020a). However, since a perfect blockage correction would be expected to collapse the $C_{P}$ and $C_{T}$ measured at different blockage ratios (e.g., figure 6) onto common unconfined $C_{P}-\lambda$ and $C_{T}-\lambda$ curves, such corrections imply not just a connection between confined and unconfined performance, but rather a broader self-similarity in turbine performance across blockage ratios. Here, we explore such self-similar aspects of array performance with a focus on the significance of the bypass flow.
To investigate the relationship between the bypass flow and array performance, we apply a method introduced by Whelan et al. (2009) and based on the bluff-body theory of Maskell (1963) to scale the experimental array performance by the corresponding bypass velocities predicted from LMADT:
@@
As shown in figure 17, the resulting ${C_{P}}^{b}-{\lambda}^{b}$ and ${C_{T}}^{b}-{\lambda}^{b}$ curves collapse across the tested blockage and tip-speed ratios, indicating that array power, thrust, and their relationship to rotation rate scale with the bypass velocity for these experimental conditions. Consequently, one may interpret $u_{b}/U_{\infty}$ as a similarity variable for array performance that is well-predicted by linear momentum theory. Some spread in the bypass-scaled thrust curves is observed (maximum $\Delta{C_{T}}^{b}\approx 0.06$), particularly at lower $\lambda^{b}$, but the bypass-scaled power curves exhibit excellent collapse at all $\lambda^{b}$ (maximum $\Delta{C_{P}}^{b}\approx 0.016$). As $u_{b}$ is calculated using the measured $C_{T}$ in Eqs. 8 and 9, it is not surprising that $C_{T}$ scales with $u_{b}$. However, given that measurements of $C_{P}$ are not involved in the calculation of $u_{b}$, the collapse in the ${C_{P}}^{b}-{\lambda}^{b}$ curves is notable, especially since the rotation, torque, and lateral forces experienced by the cross-flow turbines are all neglected in LMADT.
@@
The hydrodynamics that underlie the connection between the bypass velocity and array performance are well-described by Nishino & Willden (2012a). In summary, as $\beta$ increases, the velocity of the fluid bypassing the array increases (figures 9 and 12). Since energy is conserved in the bypass flow, this acceleration corresponds to an increased pressure drop in the bypass. To satisfy pressure equilibrium between the core and bypass flows both far upstream and far downstream, the pressure drop in the core flow through the turbines must also increase. In an open-channel flow, this pressure drop manifests as a free surface drop along the channel. As $\beta$ increases, the increased pressure drop across the array accelerates the flow through the turbines, which increases power and thrust at higher blockage.
@@
This connection between the bypass velocity and array performance parallels the theory of Maskell (1963), who proposed that the drag force on a bluff body in confined flow is related to the flow bypassing the body, not the undisturbed upstream condition. Assuming a uniform and unidirectional inflow and axisymmetric wake, Maskell developed a model for this behavior by combining conservation of momentum in the flow outside the wake with an empirical relationship for wake contraction due to confinement, and demonstrated through experiments that this model was descriptive of the drag on square flat plates at relatively low blockage ratios ($\beta\leq 10\%$). Maskell leveraged this relationship to propose a blockage correction that predicted the drag on a bluff body in unconfined flow given thrust measurements on the same body in a wind tunnel. As the wake characteristics of wind and water turbines resemble those of a bluff body under certain conditions (Medici & Alfredsson, 2006; Chamorro et al., 2012; Araya et al., 2017), several subsequent studies have extended Maskell’s theory. While some have directly applied Maskell’s flat plate blockage correction to turbine performance data, (Zilic de Arcos et al., 2020; Zhang et al., 2023), others have adapted this theory to specific turbine designs through analogous empirical relationships (Alexander & Holownia, 1978; Ross & Altman, 2011; Jeong et al., 2018; Zhang et al., 2023).
@@
Whelan et al. (2009) were the first to combine the core idea of Maskell’s theory—that the thrust on a bluff body in confined flow scales with the bypass flow—with LMADT to develop a blockage correction for “highly” loaded turbines. Rather than directly implementing Maskell’s blockage correction, Whelan et al. utilized an open-channel LMADT model to calculate the expected $u_{b}$ from axial-flow turbine thrust measurements at high blockage ($\beta=64\%$), and utilized Eqs. 17 and 18 to scale the measured thrust coefficients and tip-speed ratios by this bypass velocity. The resulting ${C_{T}}^{b}-{\lambda}^{b}$ data agreed closely with the unconfined thrust coefficients predicted for the same turbine simulated using blade element momentum (BEM) theory. Kinsey & Dumas (2017) applied the same correction to CFD-simulated thrust data for a cross-flow turbine at $\beta=13\%$, $25\%$, and $51\%$ with $u_{b}$ calculated via closed-channel LMADT. The resulting ${C_{T}}^{b}-\lambda^{b}$ curves yielded better agreement with the simulated unconfined thrust coefficients than was obtained with the popular Barnsley & Wellicome (1990) blockage correction. While Whelan et al. left open the question of whether power could be similarly corrected, Ross & Polagye (2020a) introduced Eq. 16, used the LMADT model of Houlsby et al. (2008) to calculate $u_{b}$, and scaled the corresponding performance data for both an axial-flow turbine ($\beta=2\%$ and $35\%$) and cross-flow turbine ($\beta=3\%$ and $36\%$). However, in contrast to Whelan et al. and Kinsey & Dumas, at tip-speed ratios beyond optimal ($\lambda>5$ for axial-flow, $\lambda>1.85$ for cross-flow) Ross & Polagye observed moderately poor agreement between the measured low-blockage turbine performance and ${C_{P}}^{b}$ and ${C_{T}}^{b}$ calculated from high-blockage turbine performance. Relative to prior work, the bypass-scaled power and thrust coefficients in figure 17 exhibit the strongest collapse across blockages observed to date, and highlight the similarities between array hydrodynamics and bluff-body dynamics for a wide range of high-blockage operating conditions.
@@
Since the scalings in Eqs. 16, 17 and 18 are typically interpreted as a Maskell-inspired bluff-body blockage correction, it is relevant to compare the bypass-scaled performance (figure 17) to that obtained via a “standard” LMADT blockage correction based on the theory of Glauert (1935), in which it is assumed that the thrust on the turbine responds to the velocity through the turbine ($u_{t}$). Using the LMADT-modeled values of $u_{t}$ from Eq. 10, the unconfined freestream velocity, ${U_{\infty}}^{\prime}$, that yields the same $u_{t}$ and dimensional thrust on the array in unconfined flow as in the confined system is given as
@@
The unconfined power coefficient (${C_{P}}^{\prime}$), thrust coefficient (${C_{T}}^{\prime}$), and tip-speed ratio ($\lambda^{\prime}$) are then calculated using Eqs. 16, 17 and 18 with ${U_{\infty}}^{\prime}$ in place of $u_{b}$. As shown in figure 17, up to ${C_{T}}^{\prime}\approx 1$, the ${C_{P}}^{\prime}-\lambda^{\prime}$ and ${C_{T}}^{\prime}-\lambda^{\prime}$ curves exhibit similar collapse across blockages to bluff-body scaling (maximum $\Delta{C_{P}}^{\prime}\approx 0.018$; maximum $\Delta{C_{T}}^{\prime}\approx 0.04$), with the values of ${C_{P}}^{\prime}$ and ${C_{T}}^{\prime}$ generally exceeding those of ${C_{P}}^{b}$ and ${C_{T}}^{b}$ at comparable re-scaled tip-speed. However, for LMADT in unconfined flow, the thrust coefficient cannot exceed unity as this implies reversed flow in the turbine wake. Thus, above the $\lambda^{\prime}$ corresponding to ${C_{T}}^{\prime}\approx 1$, only non-physical solutions exist, and the unconfined efficiency and thrust coefficient curves no longer collapse. Given this constraint on ${C_{T}}^{\prime}$, a key advantage of Maskell-inspired bluff-body scaling over a conventional LMADT blockage correction is the ability to describe the performance of high-blockage turbines over a wider range of confined $C_{T}$, and thus a wider range of $\beta$ and $\lambda$.
@@
In the present work, measurements of array performance at $\beta\approx 0\%$ are not available, such that the accuracy of Eqs. 16, 17 and 18 for predicting unconfined performance cannot be assessed directly. However, we expect this Maskell-inspired scaling relationship to describe self-similar array performance only under conditions where turbine dynamics resemble those of a bluff body. While the results in this study demonstrate this is likely the case for $\beta=30-55\%$, Araya et al. (2017) showed that cross-flow turbine wakes at even lower blockage ($\beta=20\%$) resemble those of a solid cylinder as both the tip-speed ratio and rotor solidity increase. Consequently, we hypothesize that the range of blockage ratios over which bluff-body scaling can capture self-similar aspects of array performance depends on operating condition and rotor geometry. Based on the results of Araya et al. (2017), we expect the lower threshold of $\beta$ at which bluff-body scaling is applicable may depend on the “dynamic solidity” of the turbine or array, which increases with increasing tip-speed ratio and rotor solidity. This may explain why the ${C_{T}}^{b}-\lambda^{b}$ curves in figure 17 increasingly collapse as $\lambda$ (and thus, $C_{T}$) increases, and is consistent with Whelan et al.’s (2009) original framing of Eqs. 17 and 18 as a blockage correction for highly-loaded turbines. Similarly, the poorer agreement between bluff-body scaled performance at $\beta=36\%$ and low-blockage performance at $\beta=3\%$ observed by Ross & Polagye (2020a) (who used a cross-flow turbine with similar solidity to this study) suggests limitations of bluff body scaling for describing low-blockage or unconfined performance.
5 Prediction of Confined Performance
@@
The self-similar array performance yielded by Maskell-inspired bluff-body scaling provides a pathway for predicting performance at different blockages from measurements or simulations at another. For example, such information would be useful for describing time-varying performance in a tidal channel with changing water depth and, therefore, blockage. Although “correcting” performance to non-zero blockages using LMADT is briefly mentioned in prior work (e.g., Steiros et al., 2022; Dehtyriov et al., 2023), there have been limited demonstrations of such methods and we are unaware of any details describing the implementation. Whelan et al. (2009) applied their Maskell-inspired bluff-body blockage correction in reverse to predict the thrust coefficients of a turbine at $\beta=64\%$ from unconfined thrust calculated from BEM, but did not detail the procedure or quantify prediction accuracy. Kinsey & Dumas (2017) proposed an empirical forecasting method based on their observation that both the power and thrust coefficients exhibit approximately linear relationships with $\beta$ at constant $\lambda$ (similar to that in figure 7). Specifically, at several $\lambda$, Kinsey & Dumas develop a linear regression for confined performance from simulations at $\beta=20\%$ and the corresponding performance at $\beta=0\%$ estimated using the LMADT blockage correction of Barnsley & Wellicome (1990). Performance at other $\beta$ was then obtained by interpolation or extrapolation. Although this method was able to accurately predict the power and thrust coefficients at $\beta=10\%$, predictions were less accurate for extrapolation to $\beta\geq 50\%$ where the $C_{T}-\beta$ and $C_{P}-\beta$ relationships are expected to be less linear.
@@
Here, a method for forecasting confined performance based on Houlsby et al.’s LMADT model and the Maskell-inspired bluff-body relationship between turbine performance and the bypass velocity is described. We expect that this approach is likely similar to that employed by Whelan et al. (2009). Consider a turbine in a channel (or an array of equally-spaced turbines spanning the channel width) with known performance at blockage ratio $\beta_{1}$ and Froude number $Fr_{h}$. At a given operating condition (e.g., $\lambda_{1}$) the turbine’s thrust coefficient is $C_{T,1}$ and bypass velocity is $u_{b,1}$. As for Glauert-derived blockage corrections, for the same turbine or array operating at a different blockage ratio, $\beta_{2}$, assume that there is a freestream velocity, $U_{\infty,2}$, that yields the same dimensional thrust on the turbine as at $\beta_{1}$ and $U_{\infty,1}$:
@@
In Eqs. 21 and 24, it is implicitly assumed that $\rho_{1}$ = $\rho_{2}$ (and thus, $\nu_{1}$ = $\nu_{2}$), which implies a difference in the freestream $Re_{D}$ (Eq. 4) between $\beta_{1}$ and $\beta_{2}$. However, the Reynolds number based on the bypass velocity ($u_{b}D/\nu$) remains constant. We note that the inability of LMADT to describe $u_{w}$ and, presumably, $u_{t}$ for relatively high confinement is not a limitation here, since LMADT is able accurately model the relationship between $\beta$, $C_{T}$, and $u_{b}/U_{\infty}$.
@@
Using $u_{b}$ as a bridge between $\beta_{1}$ and $\beta_{2}$, LMADT can be used to calculate $U_{\infty,2}$ from known quantities at $\beta_{1}$. The procedure for calculating $U_{\infty,2}$ using the LMADT model of Houlsby et al. (2008) is outlined in figure 18. Here, we assume that the Froude numbers at $\beta_{1}$ and $\beta_{2}$ are constant and subcritical as for the experimental data, although constant Froude number is not mathematically required to utilize this method. First, $u_{b}$ is determined at $\beta_{1}$ by numerically solving Eqs. 8 and 9, using $C_{T,1}$, $U_{\infty,1}$, and $Fr_{h}$ as inputs. Next, a reasonable value of $U_{\infty,2}$ is guessed (e.g., $U_{\infty,2}=U_{\infty,1}$), from which a corresponding $C_{T,2}$ is calculated from Eq. 21. Equations 8 and 9 are then solved separately for $u_{w}$ at $\beta_{2}$ using $u_{b}$, $Fr_{h}$, and the guesses for $U_{\infty,2}$ and $C_{T,2}$. The resulting values of $u_{w}$ are compared, and this process is iterated with a new guess for $U_{\infty,2}$ until the value of $U_{\infty,2}$ that minimizes the difference between Eq. 8 and Eq. 9 is determined. With $U_{\infty,2}$ known, $C_{T,2}$, $C_{P,2}$ and $\lambda_{2}$ may be calculated from $C_{T,1}$, $C_{P,1}$, and $\lambda_{1}$ via Eqs. 21, 24 and 25. Note, however, that $C_{T,1}$ is the only performance metric at $\beta_{1}$ that influences the value of $U_{\infty,2}$.
In the limiting case of $\beta_{2}\rightarrow 0$, $U_{\infty,2}\rightarrow u_{b}$ since the freestream velocity and bypass velocity are equal in unconfined flow. Consequently, Eqs. 21, 24 and 25 revert to Eqs. 16, 17 and 18 corresponding to a Maskell-inspired blockage correction (figure 17) with the implicit assumption that a bluff-body model is still representative of turbine dynamics at low blockage. Conversely, as $\beta_{2}\rightarrow 1$, the solution space is constrained by the requirements of equal bypass velocity and $Fr_{h}<1$ at $\beta_{1}$ and $\beta_{2}$, such that physically meaningful subcritical solutions at $\beta_{2}$ do not exist for all $\beta_{1}$ and $C_{T,1}$. The maximum value of $\beta_{2}$ at which a physical solution can be obtained depends on $C_{T,1}$, $\beta_{1}$, and the Froude numbers at $\beta_{1}$ and $\beta_{2}$; a visual representation of this limit for the experimental data is provided as supplementary material. However, for the performance data in this study, physically meaningful solutions exist for all practically-achievable blockages (i.e., $\beta_{2}\leq 65\%$).

### v2

@@ IV.4 Bluff-Body Scaling of Performance Data
[…]
Figure 17: Time-averaged array performance from Figure 6 scaled by 17 the analytical bypass velocities obtained using the model of Houlsby et al. [1], 17 the unconfined freestream velocities from the model of Houlsby et al. [1], 17 the analytical bypass velocities obtained using the model of Steiros et al. [24], and 17 the unconfined freestream velocities from the model of Steiros et al. [24].
To investigate the relationship between the bypass flow and array performance, we apply a method introduced by Whelan et al. [2] and inspired by the bluff-body theory of Maskell [3] to scale the experimental array performance by the corresponding bypass velocities modeled in LMADT:
@@
As shown in Figure 17, the resulting ${C_{P}}^{b}-{\lambda}^{b}$ and ${C_{T}}^{b}-{\lambda}^{b}$ curves collapse across the tested blockage and tip-speed ratios, indicating that array power, thrust, and their relationship to rotation rate scale with the bypass velocity for these experimental conditions. Consequently, one may interpret $u_{b}/U_{\infty}$ as a similarity variable for array performance that is well-predicted by LMADT. Some residual variation in the bypass-scaled thrust curves is observed, particularly at lower $\lambda^{b}$ (maximum $\Delta{C_{T}}^{b}\approx 0.06$), but the bypass-scaled power curves exhibit excellent collapse at all $\lambda^{b}$ (maximum $\Delta{C_{P}}^{b}\approx 0.016$). As $u_{b}$ is calculated using the measured $C_{T}$ in Eqs. 8 and 9, it is not surprising that $C_{T}$ scales with $u_{b}$. However, given that measurements of $C_{P}$ are not involved in the calculation of $u_{b}$, the collapse in the ${C_{P}}^{b}-{\lambda}^{b}$ curves is notable, especially since the rotation, torque, and lateral forces experienced by the cross-flow turbines are all neglected in LMADT.
@@
The hydrodynamics that underlie the connection between the bypass velocity and array performance are well-described by Nishino and Willden [10]. To summarize their conclusions, as $\beta$ increases, the velocity of the fluid bypassing the array increases (Figures 9 and 12). Since energy is mostly conserved in the bypass flow, this requires an increased pressure drop in the bypass. To satisfy pressure equilibrium between the core and bypass flows both far upstream and far downstream, the pressure drop in the core flow through the turbines must also increase. In an open-channel flow, this pressure drop manifests as a free surface drop along the channel. As $\beta$ increases, the increased pressure drop across the array accelerates the flow through the turbines, which increases power and thrust at higher blockage.
@@
This connection between the bypass velocity and array performance parallels the theory of Maskell [3], who proposed that the drag force on a bluff body in confined flow is related to the flow bypassing the body, not the undisturbed upstream condition. Assuming a uniform and unidirectional inflow and axisymmetric wake, Maskell developed a model for this behavior by combining conservation of momentum in the flow outside the wake with an empirical relationship for wake contraction due to confinement, and demonstrated through experiments that this model was descriptive of the drag on square flat plates at relatively low blockage ratios ($\beta\leq 10\%$). Maskell leveraged this relationship to propose a blockage correction that predicted the drag on a bluff body in unconfined flow given thrust measurements on the same body in a wind tunnel. As the wake characteristics of wind and water turbines resemble those of a bluff body under certain conditions [68, 69, 70], several subsequent studies have extended Maskell’s theory. While some have directly applied Maskell’s flat plate blockage correction to turbine performance data, [25, 26], others have adapted this theory to specific turbine designs through analogous empirical relationships [71, 21, 22, 26].
@@
Whelan et al. [2] were the first to combine the core idea of Maskell’s theory—that the thrust on a bluff body in confined flow scales with the bypass flow—with LMADT to develop a blockage correction for “highly loaded” turbines. Rather than directly implementing Maskell’s blockage correction, Whelan et al. utilized an open-channel LMADT model to calculate the expected $u_{b}$ from axial-flow turbine thrust measurements at high blockage ($\beta=64\%$), and utilized Eqs. 17 and 18 to scale the measured thrust coefficients and tip-speed ratios by this bypass velocity. The resulting ${C_{T}}^{b}-{\lambda}^{b}$ data agreed closely with the unconfined thrust coefficients predicted for the same turbine simulated using blade element momentum theory. Kinsey and Dumas [17] applied the same correction to CFD-simulated thrust data for a cross-flow turbine at $\beta=13\%$, $25\%$, and $51\%$ with $u_{b}$ calculated via closed-channel LMADT. The resulting ${C_{T}}^{b}-\lambda^{b}$ curves yielded better agreement with the simulated unconfined thrust coefficients than was obtained with the popular Barnsley and Wellicome [28] blockage correction. While Whelan et al. left open the question of whether power could be similarly corrected, Ross and Polagye [23] introduced Eq. 16 as an analogous scaling for $C_{P}$. Using the LMADT model of Houlsby et al. [1] to calculate $u_{b}$, Ross and Polagye scaled the corresponding performance data for both an axial-flow turbine ($\beta=2\%$ and $35\%$) and cross-flow turbine ($\beta=3\%$ and $36\%$). However, in contrast to Whelan et al. and Kinsey and Dumas, at tip-speed ratios beyond optimal ($\lambda>5$ for axial-flow, $\lambda>1.85$ for cross-flow) Ross and Polagye observed moderately poor agreement between the measured low-blockage turbine performance and ${C_{P}}^{b}$ and ${C_{T}}^{b}$ calculated from high-blockage turbine performance. Relative to prior work, the bypass-scaled power and thrust coefficients in Figure 17 exhibit the strongest collapse across blockages observed to date, and highlight similarities between array hydrodynamics and bluff-body dynamics for a wide range of high-blockage operating conditions.
@@
Since the scalings in Eqs. 16, 17 and 18 are typically interpreted as a Maskell-inspired bluff-body blockage correction, it is relevant to compare the bypass-scaled performance (Figure 17) to that obtained via a “standard” LMADT blockage correction based on the theory of Glauert [27], which assumes that the thrust on the turbine responds to the velocity through the turbine ($u_{t}$), not the bypass velocity ($u_{b}$). Using the LMADT-modeled values of $u_{t}$ from Eq. 10, the unconfined freestream velocity, ${U_{\infty}}^{\prime}$, that corresponds to the same $u_{t}$ and dimensional thrust in unconfined flow as for the confined system is given as
@@
The unconfined power coefficient (${C_{P}}^{\prime}$), thrust coefficient (${C_{T}}^{\prime}$), and tip-speed ratio ($\lambda^{\prime}$) are then calculated using Eqs. 16, 17 and 18 with ${U_{\infty}}^{\prime}$ in place of $u_{b}$. As shown in Figure 17, up to ${C_{T}}^{\prime}\approx 1$, the ${C_{P}}^{\prime}-\lambda^{\prime}$ and ${C_{T}}^{\prime}-\lambda^{\prime}$ curves exhibit similar collapse across blockages to bluff-body scaling (maximum $\Delta{C_{P}}^{\prime}\approx 0.018$; maximum $\Delta{C_{T}}^{\prime}\approx 0.04$), with the values of ${C_{P}}^{\prime}$ and ${C_{T}}^{\prime}$ generally exceeding those of ${C_{P}}^{b}$ and ${C_{T}}^{b}$ at comparable re-scaled tip-speed. However, for LMADT in unconfined flow, the thrust coefficient cannot exceed unity as this requires reversed flow in the turbine wake. Thus, above the $\lambda^{\prime}$ corresponding to ${C_{T}}^{\prime}\approx 1$, only non-physical solutions exist, and the unconfined efficiency and thrust coefficient curves no longer collapse. Given this constraint on ${C_{T}}^{\prime}$, a key advantage of Maskell-inspired bluff-body scaling over a conventional Glauert-derived scaling is the ability to describe the performance of high-blockage turbines over a wider range of confined $C_{T}$—and thus a wider range of $\beta$ and $\lambda$—since this scaling does not rely on the relationship between $C_{T}$ and $u_{w}$ in unconfined flow.
Since the potential flow model proposed by Steiros et al. [24] also relaxes the constraint on ${C_{T}}^{\prime}$, the question naturally arises as to whether this model—either coupled with Glauert’s theory as in Steiros et al. [24] or with an adaptation of Maskell’s theory following Whelan et al. [2]—can provide a better prediction of flow physics and performance self-similarity for these experiments. As shown in Figure 17, Steiros et al.’s model improves on the weaknesses of traditional blockage corrections at high $\lambda^{\prime}$ while providing similar ${C_{P}}^{\prime}$ and ${C_{T}}^{\prime}$ to that obtained with Houlsby et al.’s model at lower $\lambda^{\prime}$. Furthermore, a Maskell-inspired scaling using the $u_{b}$ calculated from Steiros et al.’s model (Figure 17) achieves a comparable collapse to that based on Houlsby et al.’s model for ${C_{T}}^{b}$ (albeit collapsing to higher values), but somewhat worse collapse for ${C_{P}}^{b}$, particularly at the highest blockages tested. Overall, the scalings in Figure 17 demonstrate the descriptiveness of Maskell-inspired bluff-body scaling for array performance is not exclusive to a particular analytical model type. Notably, despite neglecting free surface deformation, Steiros et al.’s model provides slightly more accurate estimates of the experimental bypass velocities, although the measured wake velocity remains poorly predicted by this model (see supplemental information [59]). However, because Steiros et al.’s model does not provide a materially better scaling for performance, we continue with the open-channel LMADT model of Houlsby et al. for the remainder of this study.
@@
In the present work, measurements of array performance at $\beta\approx 0\%$ are not available, such that the accuracy of Eqs. 16, 17 and 18 for predicting unconfined performance cannot be assessed directly. However, we expect this Maskell-inspired scaling relationship to describe self-similar array performance only under conditions where turbine dynamics resemble those of a bluff body. While the results in this study demonstrate this is likely the case for $\beta=30-55\%$, Araya et al. [70] showed that cross-flow turbine wakes at even lower blockage ($\beta=20\%$) resemble those of a solid cylinder as the “dynamic solidity” of the turbine increases via increasing tip-speed ratio or rotor solidity. Consequently, we hypothesize that the range of blockage ratios over which bluff-body scaling can capture self-similar aspects of array performance depends on the dynamic solidity of the turbine or array. This may explain why the ${C_{T}}^{b}-\lambda^{b}$ curves in Figures 17 and 17 increasingly collapse as $\lambda$ (and thus, $C_{T}$) increases, and is consistent with Whelan et al.’s [2] original framing of Eqs. 17 and 18 as a blockage correction for highly-loaded turbines. Similarly, the poorer agreement between bluff-body scaled performance at $\beta=36\%$ and low-blockage performance at $\beta=3\%$ observed by Ross and Polagye [23] (who used a cross-flow turbine with similar solidity to this study) suggests limitations of bluff body scaling for describing low-blockage or unconfined performance. In these regimes, the blockage correction proposed by Steiros et al. [24] (Figure 17) has clear advantages for high-solidity turbines.
V Prediction of Confined Performance
@@
The self-similar array performance yielded by Maskell-inspired bluff-body scaling provides a pathway for predicting performance at a given blockage ratio using measurements or simulations at another. For example, such information would be useful for describing time-varying performance in a tidal channel with changing water depth and, therefore, blockage. However, only a handful of prior studies have considered “correcting” performance to non-zero blockages. Whelan et al. [2] applied their Maskell-inspired bluff-body blockage correction in reverse to predict the thrust coefficients of a turbine at $\beta=64\%$ from unconfined thrust calculated from BEM, but did not detail the procedure or quantify prediction accuracy. Kinsey and Dumas [17] proposed an empirical forecasting method based on their observation that both the power and thrust coefficients exhibit approximately linear relationships with $\beta$ at constant $\lambda$ (similar to that in Figure 7). Specifically, at several $\lambda$, Kinsey and Dumas develop a linear regression for confined performance from simulations at $\beta=20\%$, and estimated the corresponding performance at $\beta=0\%$ using the LMADT blockage correction of Barnsley and Wellicome [28]. Performance at other $\beta$ was then obtained by interpolation or extrapolation. Although this method was able to accurately predict the power and thrust coefficients at $\beta=10\%$, predictions were less accurate for extrapolation to $\beta\geq 50\%$ where the $C_{T}-\beta$ and $C_{P}-\beta$ relationships are less linear. Finally, the blockage correction developed by Steiros et al. [24] can be used to “correct” turbine performance to arbitrary blockage by simply specifying a non-zero target value of $\beta$ in their equation set. While Steiros et al. explicitly highlight this capability, they do not demonstrate it or evaluate prediction accuracy.
@@
Here, a method for forecasting confined performance based on Houlsby et al.’s LMADT model and the Maskell-inspired bluff-body scaling is described. We expect that this approach is likely similar to that employed by Whelan et al. [2]. A separate evaluation of the forecasting method of Steiros et al. [24] as applied to the experimental data is provided as supplementary material [59]. Consider a turbine in a channel (or an array of equally-spaced turbines spanning the channel width) with known performance at blockage ratio $\beta_{1}$ and Froude number $Fr_{h}$. At a given operating condition (e.g., $\lambda_{1}$) the turbine’s thrust coefficient is $C_{T,1}$ and bypass velocity is $u_{b,1}$. As for Glauert-derived blockage corrections, for the same turbine or array operating at a different blockage ratio, $\beta_{2}$, assume that there is a freestream velocity, $U_{\infty,2}$, that yields the same dimensional thrust on the turbine as at $\beta_{1}$ and $U_{\infty,1}$:
@@
In Eqs. 21 and 24, it is implicitly assumed that $\rho_{1}$ = $\rho_{2}$ (and thus, $\nu_{1}$ = $\nu_{2}$), which implies a difference in the freestream $Re_{D}$ (Eq. 4) between $\beta_{1}$ and $\beta_{2}$. However, the Reynolds number based on the bypass velocity ($u_{b}D/\nu$) remains constant. We note that the inability of LMADT to describe $u_{w}$ and, presumably, $u_{t}$ for relatively high confinement is not a limitation here, since LMADT is able to accurately model the relationship between $\beta$, $C_{T}$, and $u_{b}/U_{\infty}$.
@@
Using $u_{b}$ as a bridge between $\beta_{1}$ and $\beta_{2}$, LMADT can be used to calculate $U_{\infty,2}$ from known quantities at $\beta_{1}$. The procedure for calculating $U_{\infty,2}$ using the LMADT model of Houlsby et al. [1] is outlined in Figure 18. Here, we assume that the Froude numbers at $\beta_{1}$ and $\beta_{2}$ are constant and subcritical as for the experimental data, although constant Froude number is not required. First, $u_{b}$ is determined at $\beta_{1}$ by numerically solving Eqs. 8 and 9, using $C_{T,1}$, $U_{\infty,1}$, and $Fr_{h}$ as inputs. Next, a reasonable value of $U_{\infty,2}$ is guessed (e.g., $U_{\infty,2}=U_{\infty,1}$), from which a corresponding $C_{T,2}$ is calculated from Eq. 21. Equations 8 and 9 are then solved separately for $u_{w}$ at $\beta_{2}$ using $u_{b}$, $Fr_{h}$, and the guesses for $U_{\infty,2}$ and $C_{T,2}$. The resulting values of $u_{w}$ are compared, and this process is iterated with a new guess for $U_{\infty,2}$ until the value of $U_{\infty,2}$ minimizes the difference between Eq. 8 and Eq. 9. With $U_{\infty,2}$ known, $C_{T,2}$, $C_{P,2}$ and $\lambda_{2}$ may be calculated from $C_{T,1}$, $C_{P,1}$, and $\lambda_{1}$ via Eqs. 21, 24 and 25. Note, however, that $C_{T,1}$ is the only performance metric at $\beta_{1}$ that influences the value of $U_{\infty,2}$.
In the limiting case of $\beta_{2}\rightarrow 0$, $U_{\infty,2}\rightarrow u_{b}$ since the freestream velocity and bypass velocity are equal in unconfined flow. Consequently, Eqs. 21, 24 and 25 revert to Eqs. 16, 17 and 18 corresponding to a Maskell-inspired blockage correction (Figure 17) with the implicit assumption that a bluff-body model is still representative of turbine dynamics at low blockage. Conversely, as $\beta_{2}\rightarrow 1$, the solution space is constrained by the requirements of equal bypass velocity and $Fr_{h}<1$ at $\beta_{1}$ and $\beta_{2}$, such that physically meaningful subcritical solutions at $\beta_{2}$ do not exist for all $\beta_{1}$ and $C_{T,1}$. The maximum value of $\beta_{2}$ at which a physical solution can be obtained depends on $C_{T,1}$, $\beta_{1}$, and the Froude numbers at $\beta_{1}$ and $\beta_{2}$; a visual representation of this limit for the experimental data is provided as supplementary material [59]. However, for the performance data in this study, physically meaningful solutions exist for all practically-achievable blockages (i.e., $\beta_{2}\leq 65\%$).